### v3

# Modeling Spatio-Temporal Transport: From Rigid Advection to Realistic Dynamics

Abstract
Stochastic models for spatio-temporal transport face a critical trade-off between physical realism and interpretability. The advection model with a single constant velocity is interpretable but physically limited by its perfect correlation over time. This work aims to bridge the gap between this simple framework and its physically realistic extensions. Our guiding principle is to introduce a spatial correlation structure that vanishes over time. To achieve this, we present two distinct approaches. The first constructs complex velocity structures, either through superpositions of advection components or by allowing the velocity to vary locally. The second is a spectral technique that replaces the singular spectrum of rigid advection with a more flexible form, introducing temporal decorrelation controlled by parameters. We accompany these models with efficient simulation algorithms and demonstrate their success in replicating complex dynamics, such as tropical cyclones and the solutions of partial differential equations.
Keywords: Spatio-temporal models, Non-separable models, Advection-diffusion processes, Simulation, Velocity fields, Spectral methods
1 Introduction
Spatio-temporal phenomena driven by advection are fundamental to a vast array of scientific disciplines, from tracking atmospheric pollutants and forecasting ocean currents to modeling seismic wave propagation. The modeling of these processes is challenging due to their complex interdependent evolution in space and time. While the field of spatio-temporal statistics has made substantial progress, with seminal works like Cressie2011, accurately capturing the dynamics of transport remains an active and critical area of research.
To model these dynamics, one major research direction uses Lagrangian approaches to build non-stationary covariance structures by tracking particle trajectories. This framework has proven effective for modeling environmental processes like ocean waves and precipitation (ailliot2011; benoit2018), as these phenomena are inherently governed by physical transport. Another line of research employs Stochastic Partial Differential Equations (SPDEs) to embed physical principles like advection and diffusion directly into the models. This has led to computationally efficient models with interpretable parameters (sigrist2015), which have since been extended to handle more complex, spatially-varying dynamics (liu2022; carlotto2024). While other methods focus on simulating broader classes of non-separable covariance functions (see for instance allard2020simulating), they often face a trade-off: they either omit explicit directional transport for simplicity, or incorporate it at the cost of significant model complexity. Our work seeks to occupy a middle ground, prioritizing both physical realism and interpretability. For comprehensive reviews on the state of dynamic and non-separable models, a topic that builds on foundational work like gneiting2002nonseparable, porcu2020nonseparable, chen2021space, and papalexiou2021advancing.
To achieve physical realism, our work begins with the foundational model of a spatial random field advected by a constant velocity—a classic benchmark used to approximate natural phenomena in works like cox1988. Our goal is to develop extensions from this starting point that add complexity while retaining the original’s intuitive interpretability. We refer to the simplest motivating model as the frozen field, following the terminology of christakos2017. The advantage of this starting point is its structure. In particular, its value, covariance, and spectrum are all defined by three simple components: the initial spatial field, its covariance function, and the constant velocity. This allows us to systematically manipulate these components to build more complex and realistic models while maintaining the clear physical interpretation of the original.
To this end, we explore two distinct strategies to generalize the frozen field. The first strategy focuses on enriching the velocity structure itself, moving beyond a single constant vector. This leads to two models, the first of which, related to earlier work on sums of advected fields (schlather2010some), relies on the superposition of multiple frozen fields, while the second considers a velocity that is itself a vector field varying in space and time. Our second strategy involves relaxing the spectral shape of the frozen field, replacing its singular frequency-domain representation with more flexible forms that allow for natural decorrelation over time, which can be controlled with parameters in the model.
To complement our theoretical analysis, we also develop efficient simulation methods, which are vital for model validation, parameter inference, and prediction. Our approach to simulation mirrors our theoretical strategy: we first establish methods for the frozen field model, and then extend them to more complex frameworks. Specifically, we begin by introducing two distinct procedures to simulate a frozen field, one in the space-time domain and the other in the spectral domain. We then demonstrate how these algorithms can be adapted to simulate the novel models proposed in this work.
Thus, this paper makes a threefold contribution. First, we introduce these novel classes of models for advection. Second, we provide computationally efficient algorithms to simulate them, ensuring that these methods are accessible for practical application. We demonstrate that these algorithms are competitive with, and in some cases superior to, standard approaches by avoiding known issues like numerical diffusion. Third, we demonstrate the applicability of these models by demonstrating their ability to generate complex, physically plausible dynamics, such as those of a tropical cyclone, and to replicate the solutions of well-known transport and diffusion PDEs. To facilitate their use by practitioners, we provide implementations of these methods in R (Rmanual).
The paper is organized as follows. Section 2 formally defines the frozen field model and discusses two primary simulation approaches. Section 3 introduces our first class of generalizations based on velocity structures, presents simulated examples, and compares the results to numerical PDE solvers. Section 4 details our second generalization strategy based on spectral modifications and analyzes the role of the model parameters. Finally, Section 5 provides concluding remarks and outlines directions for future research. Further results are included in the supplementary materials.
2 Space Random Fields with Advection
In what follows, we give a formal definition of a spatial random field with advection, including the analytical formula of its covariance and spectral density functions. Then, we propose two numerical algorithms to efficiently simulate this class of phenomena.
2.1 Framework and definition
We begin by defining fundamental concepts for spatial random fields, which form the basis for the spatio-temporal models discussed later. Consider a continuous-space random field $X_{S}:\mathbb{R}^{d}\to\mathbb{R}$, where $X_{S}(\mathbf{s})$ is the value of the field at the spatial coordinate $\mathbf{s}\in\mathbb{R}^{d}$. For simplicity, we consider $d=2$, though extensions to higher dimensions are generally straightforward. Assuming $X_{S}(\mathbf{s})$ is a Gaussian random field, its probabilistic structure is fully characterized by its mean function $\mathrm{E}\{X_{S}(\mathbf{s})\}$ and its covariance function $c_{XX}(\mathbf{s}_{1},\mathbf{s}_{2})=\mathrm{cov}\{X_{S}(\mathbf{s}_{1}),X_{S}(\mathbf{s}_{2})\}$ for any spatial locations $\mathbf{s},\mathbf{s}_{1},\mathbf{s}_{2}\in\mathbb{R}^{2}$.
A common and important simplifying assumption is spatial stationarity. A spatial random field $X_{S}(\mathbf{s})$ is said to be stationary if its mean is constant (assumed to be zero without loss of generality for covariance modeling) and its covariance function $c_{XX}(\mathbf{s}_{1},\mathbf{s}_{2})$ depends only on the lag $\mathbf{h}=|\mathbf{s}_{2}-\mathbf{s}_{1}|\in\mathbb{R}^{2}$. With a slight abuse of notation, we write this stationary covariance function as
$c_{XX}(\mathbf{h})=\mathrm{cov}\{X_{S}(\mathbf{s}),X_{S}(\mathbf{s}+\mathbf{h})\}.$
Under stationarity, the spectral analysis of $X_{S}(\mathbf{s})$ becomes particularly useful. The $2$-dimensional Fourier transform of the stationary covariance function $c_{XX}(\mathbf{h})$ yields the spectral density function (or spectrum), $S_{XX}(\mathbf{k})$:
$S_{XX}(\mathbf{k})=\int_{\mathbb{R}^{2}}c_{XX}(\mathbf{h})e^{-2\pi i\mathbf{k}^{\top}\mathbf{h}}\mathrm{d}\mathbf{h},$
where $\mathbf{k}\in\mathbb{R}^{2}$ is the wavenumber. Being the Fourier transform of the covariance function, the spectrum $S_{XX}(\mathbf{k})$ describes the contribution of different frequencies to the field’s variance.
Having established the properties of spatial random fields, we now extend these concepts to the space-time domain to model evolving phenomena. A space-time random field $Z:\mathbb{R}^{2}\times\mathbb{R}^{+}\to\mathbb{R}$ takes values over space-time coordinates $(\mathbf{s},t)\in\mathbb{R}^{2}\times\mathbb{R}^{+}$. We consider a special class of space-time random fields, namely that of traveling random fields. One can think of them as space random fields advected over time. The simplest form for such phenomena consists in a spatial field travelling with constant velocity $\mathbf{v}\in\mathbb{R}^{2}$. This entails that the random field is “frozen” and rigidly shifted in time according to the velocity. We adopt the name frozen field from christakos2017spatiotemporal to denote such space-time process. We recall the definition of this stochastic process.
Definition 2.1 (Frozen Field).
A random field $Z(\mathbf{s},t)$ is a Frozen Field if there is a constant velocity $\mathbf{v}\in{\mathbb{R}}^{2}$, and a random field $X_{S}(\mathbf{s})$ with $\mathbf{s}\in{\mathbb{R}}^{2}$ and covariance function $c_{XX}(\mathbf{h})$ such that
$Z(\mathbf{s},t)=X_{S}(\mathbf{s}-\mathbf{v}t),$ (1)
with $t\in\mathbb{R}^{+}$.
The Frozen Field (FF from here onward) $Z(\mathbf{s},t)$ is the solution of transport equation
$\begin{cases}\frac{\partial Z(\mathbf{s},t)}{\partial t}+\mathbf{v}\cdot\nabla Z(\mathbf{s},t)=0,\\ Z(\mathbf{s},0)=X_{S}(\mathbf{s}),\end{cases}$ (2)
where $\nabla Z(\mathbf{s},t)=\left(\frac{\partial Z(\mathbf{s},t)}{\partial s_{1}}\frac{\partial Z(\mathbf{s},t)}{\partial s_{2}}\right)^{\top}$. Moreover, such process enjoys the following second order properties:
$c_{ZZ}(\mathbf{h},\tau)$ $=c_{XX}(\mathbf{h}-\mathbf{v}\tau),$ (3)
$S_{ZZ}(\mathbf{k},\omega)$ $=S_{XX}(\mathbf{k})\cdot\delta(\omega+\mathbf{k}^{T}\mathbf{v}),$ (4)
where $(\mathbf{k},\omega)\in\mathbb{R}^{2}\times\mathbb{R}$ denotes the pair of wavenumber and frequency, respectively. Specifically, covariance $c_{ZZ}(\mathbf{h},\tau)$ consists of the covariance function $c_{XX}(\cdot)$ of the space random field evaluated at spatial lag $|\mathbf{h}-\mathbf{v}\tau|$, and it is clearly not separable. The spectrum $S_{ZZ}(\mathbf{k},\omega)$ depends on the spectrum of the space random field $S_{XX}(\mathbf{k})$, which is multiplied by a generalized function corresponding to a Dirac delta depending on velocity $\mathbf{v}$.
In this space-time model, spatial locations shift according to the constant velocity as time progresses. More specifically, spatial correlation $\rho(\mathbf{h},\tau)=\frac{\left|c_{XX}(\mathbf{h}-\mathbf{v}\tau)\right|}{c_{XX}(0)}$ of the FF is such that
$\rho(\mathbf{v}\tau,\tau)=\frac{c_{XX}(\mathbf{v}\tau-\mathbf{v}\tau)}{c_{XX}(0)}=1,$
namely the correlation is perfect along propagation path $(\mathbf{v}\tau,\tau)$. However, under normal circumstances, one would not expect an advection phenomenon to persist over all positive lags, as this would imply a particle traveling through a perfectly stationary medium without friction. Energy dissipation, namely the loss of kinetic energy over time, would be a more realistic expectation (christakos2017spatiotemporal). On the other hand, despite being overly-simplistic, the FF model enjoys the striking benefit of being a straightforwardly interpretable model for nonseparable space-time random fields. Relying on such interpretability and driven by the necessity for flexibility in describing complex spatiotemporal phenomena, the core of this work is to introduce new models by relaxing certain assumptions of the FF model.
As we will provide simulation procedure for the these new models, we first describe two algorithms to simulate a FF. This is the content of the next section.
2.2 Simulation of Frozen Fields
Say we wish to simulate a FF traveling with known velocity $\mathbf{v}$. Such space-time process is discretized over a spatial grid $\mathcal{G}=[0,n]\times[0,n]\in\mathbb{N}_{0}^{2}$, which is square for sake of simplicity, and time grid $\{t_{0},\dots,t_{T}\}\in\mathbb{N}_{0}$. As mentioned previously, this can be done in two ways.
Firstly, a FF can be generated by simulating a single, large spatial random field, and then advecting $\mathcal{G}$ across this field over time. More precisely, we first need to generate the extended spatial field $X_{S}^{\text{ext}}(\mathbf{s})$ on grid $\mathcal{G}^{*}=[0,N]\times[0,N]$, with $N>n$. This initial simulation can be performed with circulant embedding (ChanWood97; DietrichNewsam97; Graham2018), a non-exact spectral method relying on the use of Discrete Fourier Transforms (DFTs) and their implementation on regular grids, called Fast Fourier Transforms (FFTs). The computational cost of circulant embedding, for which a FFT is carried out once, is $\mathcal{O}(N^{2}\log N)$. Notice that the simulation domain $\mathcal{G}^{*}$ must be large enough to contain the full trajectory of the smaller observation grid $\mathcal{G}$. The required grid size $N$ should be chosen based on the properties of the spatio-temporal field being simulated. Accordingly, we set $N=n+Tv_{\text{max}}$, where $v_{\text{max}}=\max\{|v_{1}|,|v_{2}|\}$ is the maximum component of the constant velocity $\mathbf{v}=(v_{1},v_{2})^{T}$.
Once $X_{S}^{\text{ext}}$ is generated, the space-time field $Z(\mathbf{s},t)$ is constructed by sampling from this single large field. For each point $\mathbf{s}$ on the observation grid $\mathcal{G}$, and for each of the $T$ time steps, the value is given by $Z(\mathbf{s},t)=X_{S}^{\text{ext}}(\mathbf{s}-\mathbf{v}t)$. It is important to note that the advected location $\mathbf{s}-\mathbf{v}t$ will generally not coincide with a point on the discrete grid $\mathcal{G}^{*}$. Therefore, its value must be obtained through a spatial interpolation scheme, such as bilinear or bicubic interpolation. This sampling and interpolation process for all points and time steps has a total cost of $\mathcal{O}(n^{2}T)$.
A comparison of computational costs reveals a clear trade-off between this shifting method and a direct space-time simulation via 3D circulant embedding, which costs $\mathcal{O}(n^{2}T\log(nT))$. The total cost of the shifting method, $\mathcal{O}(N^{2}\log N+n^{2}T)$, depends heavily on $N$, which, by construction, grows linearly with $T$. Consequently, the shifting method is significantly more efficient when the number of time steps $T$ is small relative to the spatial grid dimension $n$. For long large $T$, the 3D spectral approach becomes more advantageous as its cost scales more gracefully with time. For context, both methods are vastly more efficient than an exact simulation using Cholesky decomposition of the full space-time covariance matrix, which would require a cost of $\mathcal{O}(n^{6}T^{3})$. The overall procedure for the shifting method is detailed in Algorithm 1.
Algorithm 1 Simulation of a Frozen Field via Grid Shifting
1:Spatial grid dimension $n$, number of time steps $T$, constant velocity $\mathbf{v}$, spatial covariance function $c_{XX}$.
2:Compute max velocity component $v_{\text{max}}\leftarrow\max\{|v_{1}|,|v_{2}|\}$
3:Compute extended grid dimension $N\leftarrow n+Tv_{\text{max}}$
4:Define extended grid $\mathcal{G}^{*}\leftarrow[0,N]\times[0,N]$ $\triangleright$ Grid that accommodates for advection
5:Generate Gaussian random field $X_{S}^{\text{ext}}$ on $\mathcal{G}^{*}$ using $c_{XX}$ $\triangleright$ Simulate large background field
6:Define simulation grid $\mathcal{G}\leftarrow[0,n]\times[0,n]$
7:Initialize space-time field $Z\in\mathbb{R}^{(n+1)\times(n+1)\times(T+1)}$
8:for each spatial location $\mathbf{s}$ in $\mathcal{G}$ do
9:  $Z(\mathbf{s},0)\leftarrow X_{S}^{\text{ext}}(\mathbf{s})$ $\triangleright$ Set initial state from the large field
10:  for $t=1$ to $T$ do
11:   $Z(\mathbf{s},t)\leftarrow X_{S}^{\text{ext}}(\mathbf{s}-\mathbf{v}t)$ $\triangleright$ Advect by sampling the large field via interpolation
12:  end for
13:end for
14:return $Z$
A second approach is an adaptation of the spectral method called Davies-Harte algorithm (see for instance craigmile2003). This method simulates the Gaussian random field based on its spectral density function. The procedure, detailed in Algorithm 2, begins by defining padded grid dimensions, $n_{\text{pad}}=2n$ and $T_{\text{pad}}=2T$. This step is crucial for mitigating the periodic boundary effects that are inherent in any FFT-based simulation.
On this padded grid, a complex-valued field, $\widetilde{Z}$, is constructed. For each point in the discrete frequency domain, the value of $\widetilde{Z}$ is set by multiplying the square root of the space-time spectrum, $\sqrt{S_{ZZ}(\mathbf{k},\omega)}$, by a random complex white noise sample. Once this complex grid is fully constructed, it is transformed to the space-time domain using a single 3D Inverse FFT (IFFT). The real part of the resulting padded field is taken, and the final $n\times n\times T$ field is obtained by cropping the result, discarding the extra grid points that were added for padding.
A key practical consideration for this spectral method is its accuracy. Since the theoretical spectrum (4) contains a Dirac delta function, which cannot be represented exactly on a discrete frequency grid, it must be approximated, for example, with a narrow Gaussian kernel. This approximation introduces a trade-off. For a given wavenumber-frequency grid resolution, a narrower kernel more accurately represents the Dirac delta, but can lead to higher numerical instability or aliasing. Conversely, a wider kernel smooths the spectrum, deviating more from the advected behavior, but offering greater numerical stability. Furthermore, as mentioned before, to mitigate artificial correlations arising from periodic boundary effects, the Davies-Harte algorithm employs padding by doubling the grid size in all dimensions. This effectively creates a buffer region where the artificial periodicities occur, ensuring that the central, desired portion of the simulated field remains largely free from such distortions. However, the overall quality of the simulation is also influenced by the discrete resolution of the frequency grid, which approximates the continuous theoretical spectrum. The computational cost of this entire procedure is dominated by the single 3D IFFT, leading to a complexity of $\mathcal{O}(n^{2}T\log(nT))$.
Algorithm 2 Simulation of a Frozen Field via its Spectrum (Davies-Harte)
1:Spatial grid dimension $n$, number of time steps $T$, space-time spectral density function $S_{ZZ}$.
2:Define padded dimensions $n_{\text{pad}}\leftarrow 2n$ and $T_{\text{pad}}\leftarrow 2T$ $\triangleright$ Double grid sizes
3:Define wavenumber vector $\mathbf{k}_{\text{vec}}$ for a grid of size $n_{\text{pad}}$
4:Define frequency vector $\omega_{\text{vec}}$ for a grid of size $T_{\text{pad}}$
5:Initialize complex grid in frequency domain $\widetilde{Z}\in\mathbb{C}^{n_{\text{pad}}\times n_{\text{pad}}\times T_{\text{pad}}}$
6:for each index $(i,j,t)$ in the padded grid dimensions do
7:  Get corresponding wavevector $\mathbf{k}\leftarrow(\mathbf{k}_{\text{vec}}[i],\mathbf{k}_{\text{vec}}[j])$ and frequency $\omega\leftarrow\omega_{\text{vec}}[t]$
8:  Evaluate spectral density $S\leftarrow S_{ZZ}(\mathbf{k},\omega)$
9:  Generate complex white noise $\gamma\sim\mathcal{CN}(0,1)$ $\triangleright$ $\Re(\gamma),\Im(\gamma)\sim\mathcal{N}(0,1)$
10:  $\widetilde{Z}[i,j,t]\leftarrow\sqrt{S}\cdot\gamma$ $\triangleright$ Construct the random field in frequency domain
11:end for
[…]
13:Take the real part $Z_{\text{pad}}\leftarrow\Re(Z_{\text{pad}})$
14:Crop to original dimensions $Z\leftarrow Z_{\text{pad}}[0:n-1,0:n-1,0:T-1]$
15:return $Z$
Having established the properties and simulation methods for the FF model, we now turn to the analysis of its extensions.
3 Generalization via Velocity Structures
This section is dedicated to the extensions of the FF model with our first strategy, namely modifying the overly simplistic velocity structure of model (1). In particular, we consider two models, one consisting of the overlapping of FFs, and one where we substitute the constant velocity with a space-time dependent vector field. We study the properties of both of them, including their propagation path.
3.1 Framework
Our first extension of model (1) consists of juxtaposition of several fields with different advection directions. To do so, consider geometric transformations of a constant velocity $\mathbf{v}\in\mathbb{R}^{2}$ in (1). Specifically, we examine composite transformations $\mathbf{M}\in\mathbb{R}^{2\times 2}$ of the form
$\mathbf{M}=\mathbf{R}\mathbf{S},$ (5)
where $\mathbf{R}\in\mathbb{R}^{2\times 2}$ is a rigid rotation around the origin and $\mathbf{S}\in\mathbb{R}^{2\times 2}$ denotes a scaling transformation.
Definition 3.1 (Distributed FF).
Consider $n_{\text{v}}$ rotations $\mathbf{R}_{1},\dots,\mathbf{R}_{n_{\text{v}}}$ and scalings $\mathbf{S}_{1},\dots,\mathbf{S}_{n_{\text{v}}}$. A random field $Z(\mathbf{s},t)$ is a Distributed Frozen Field if, given a velocity $\mathbf{v}\in\mathbb{R}^{2}$, there exists a set of transformations $\mathbf{M}_{1},\dots,\mathbf{M}_{n_{\text{v}}}$ defined as in (5), and a random field $X_{S}(\mathbf{s})$ with $\mathbf{s}\in\mathbb{R}^{2}$ and covariance function $c_{XX}(\mathbf{h})$ such that
$Z(\mathbf{s},t)=\sum_{i=1}^{n_{\text{v}}}p(\mathbf{M}_{i}\mathbf{v})Z_{i}(\mathbf{s},t)=\sum_{i=1}^{n_{\text{v}}}p(\mathbf{M}_{i}\mathbf{v})X_{S}\left(\mathbf{s}-(\mathbf{M}_{i}\mathbf{v})t\right),$ (6)
for $t\in\mathbb{R}^{+}$, where the weights satisfy
$\sum_{i=1}^{n_{\text{v}}}p(\mathbf{M}_{i}\mathbf{v})$ $=1,$
$p(\mathbf{M}_{i}\mathbf{v})$ $\geq 0,\quad\forall i=1,\dots,n_{\text{v}}.$
The constraints requiring the weights to be non-negative and sum to one are imposed to ensure the model is a well-defined convex combination of the component fields. However, while these conditions are sufficient for the model’s theoretical definition, they do not guarantee parameter identifiability in a practical estimation setting. For that purpose, additional restrictions, such as imposing an ordering on the weights, would be required. From a modeling perspective, the key advantage of this framework is its flexibility. By allowing for multiple transformations $\mathbf{M}_{i}$, the model gains flexibility in modeling various directional and scaling influences, making it suitable for applications where the propagation dynamics are heterogeneous or subject to multiple driving factors. In particular, each component of (6) is the solution of
$\begin{cases}\frac{\partial Z_{i}(\mathbf{s},t)}{\partial t}+(\mathbf{M}_{i}\mathbf{v})\cdot\nabla Z_{i}(\mathbf{s},t)=0,\\ Z_{i}(\mathbf{s},0)=X_{S}(\mathbf{s}).\end{cases}$
In order to introduce a realistic appearance to $Z(\mathbf{s},t)$, we suggest selecting the transformations $\mathbf{M}_{1},\dots,\mathbf{M}_{n_{\text{v}}}$ in (6) randomly, i.e., drawing the rotation and scaling transformations from specified probability distributions. Moreover, for each transformation $\mathbf{M}_{i}$, the weight $p(\mathbf{M}_{i}\mathbf{v})$ determines the prominence of the corresponding geometric feature in the spatio-temporal phenomenon.
Notice that considering a random velocity in equation (1) is an existing extension to the FF model. For instance, schlather2010some generates a Gaussian velocity with mean $\boldsymbol{\mu}\in\mathbb{R}^{2}$ and covariance $D\in\mathbb{R}^{2\times 2}$, i.e., $\mathbf{v}\sim\mathcal{N}(\boldsymbol{\mu},D)$, once for the entire field, while christakos2017spatiotemporal considers $\mathbf{v}^{t}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\boldsymbol{\mu},D)$, where the velocity $\mathbf{v}^{t}$ is drawn independently at every time point $t$. With the Distributed FF model, we offer a more complex alternative by allowing for random geometric transformations of the velocity, enabling multiple propagation directions and scales within the same field. The diminished correlation in the distributed case arises from the superposition of multiple transformed components introducing variability in propagation directions and scales.
Proposition 3.1.
Assume that $Z(\mathbf{s},t)$ is a distributed frozen field as specified in Definition 3.1. Then, for fixed $(\mathbf{s},t)\in\mathbb{R}^{2}\times\mathbb{R}^{+}$, the covariance function and the propagation paths of $Z(\mathbf{s},t)$ are given by
$\mathrm{cov}(Z(\mathbf{s},t),Z(\mathbf{s}+\mathbf{h},t+\tau))$ $=\sum_{i=1}^{n_{\text{v}}}p(\mathbf{M}_{i}\mathbf{v})\sum_{j=1}^{n_{\text{v}}}p(\mathbf{M}_{j}\mathbf{v})\,c_{XX}\left(\mathbf{h}-(\mathbf{M}_{j}\mathbf{v})\tau-(\mathbf{M}_{j}\mathbf{v}-\mathbf{M}_{i}\mathbf{v})t\right),$ (7)
with propagation paths
$\left((\mathbf{M}_{j}\mathbf{v})\tau+(\mathbf{M}_{j}\mathbf{v}-\mathbf{M}_{i}\mathbf{v})t,\tau\right),\quad i,j=1,\dots,n_{\text{v}}.$ (8)
Proof.
For fixed $(\mathbf{s},t)\in\mathbb{R}^{2}\times\mathbb{R}^{+}$, the covariance of $Z(\mathbf{s},t)$ and $Z(\mathbf{s}+\mathbf{h},t+\tau)$ is computed using the bilinearity of covariance:
$\mathrm{cov}(Z(\mathbf{s},t),Z(\mathbf{s}+\mathbf{h},t+\tau))$
$=\mathrm{cov}\left(\sum_{i=1}^{n_{\text{v}}}p(\mathbf{M}_{i}\mathbf{v})X_{S}\left(\mathbf{s}-(\mathbf{M}_{i}\mathbf{v})t\right),\sum_{j=1}^{n_{\text{v}}}p(\mathbf{M}_{j}\mathbf{v})X_{S}\left(\mathbf{s}+\mathbf{h}-(\mathbf{M}_{j}\mathbf{v})(t+\tau)\right)\right)$
$=\sum_{i=1}^{n_{\text{v}}}\sum_{j=1}^{n_{\text{v}}}p(\mathbf{M}_{i}\mathbf{v})p(\mathbf{M}_{j}\mathbf{v})\,\mathrm{cov}\left(X_{S}\left(\mathbf{s}-(\mathbf{M}_{i}\mathbf{v})t\right),X_{S}\left(\mathbf{s}+\mathbf{h}-(\mathbf{M}_{j}\mathbf{v})(t+\tau)\right)\right)$
$=\sum_{i=1}^{n_{\text{v}}}\sum_{j=1}^{n_{\text{v}}}p(\mathbf{M}_{i}\mathbf{v})p(\mathbf{M}_{j}\mathbf{v})\,c_{XX}\left(\mathbf{h}-(\mathbf{M}_{j}\mathbf{v})\tau-(\mathbf{M}_{j}\mathbf{v}-\mathbf{M}_{i}\mathbf{v})t\right).$
Unity correlation occurs when the argument of $c_{XX}$ is zero, i.e.,
$\mathbf{h}=(\mathbf{M}_{j}\mathbf{v})\tau+(\mathbf{M}_{j}\mathbf{v}-\mathbf{M}_{i}\mathbf{v})t,\quad i,j=1,\dots,n_{\text{v}},$
thanks to the fact that $Z(\mathbf{s},t)$ is a convex combination of frozen fields.
This proposition reveals how the Distributed FF model generalizes the classic FF by examining the structure of its covariance in (7). First, consider the terms where $i=j$. In this case, the expression simplifies to that of a standard FF: the time-dependent component $(\mathbf{M}_{j}\mathbf{v}-\mathbf{M}_{i}\mathbf{v})t$ vanishes from the covariance argument, and the propagation path (8) reduces to simple, straight-line advection, $\left((\mathbf{M}_{i}\mathbf{v})\tau,\tau\right)$. These terms represent a weighted mixture of independent FFs, each with its own transformed velocity. Conversely, the terms where $i\neq j$ introduce novel cross-covariances between the different components. Specifically, $(\mathbf{M}_{j}\mathbf{v}-\mathbf{M}_{i}\mathbf{v})t$ denotes the growing spatial displacement between two component paths over time $t$. The presence of this term makes the overall covariance function dependent on time, thus making the Distributed FF model non-stationary. This complex behavior is known in physics as interference, referring to the interactions introduced by the superimposition of waves traveling in the same medium. Consequently, the distributed frozen field satisfies $\rho(v\tau,\tau)<1$ unless $\mathbf{M}_{i}\mathbf{v}=\mathbf{v}$ for all $i$, thereby preventing unrealistic perfect correlation along the propagation path of the FF.
In some settings it is more natural to assume the random field locally satisfies the wave equation. This means that, for every spatial location $\mathbf{s}$ and time point $t$, we define a velocity $\mathbf{v}(\mathbf{s},t)$, where $\mathbf{v}:\mathbb{R}^{2}\times\mathbb{R}^{+}\to\mathbb{R}^{2}$ is a continuous vector field. With this in mind, we propose the following stochastic process:
Definition 3.2 (Evolving FF).
A random field $Z(\mathbf{s},t)$ is a Evolving Frozen Field if there is a deterministic velocity field $\mathbf{v}(\mathbf{s},t)\in{\mathbb{R}}^{2}$, and a random field $X_{S}(\mathbf{s})$ with $\mathbf{s}\in{\mathbb{R}}^{2}$ and covariance function $c_{XX}(\mathbf{h})$ such that
$Z(\mathbf{s},t)=X_{S}(\mathbf{s}-\mathbf{v}(\mathbf{s},t)t),$ (9)
with $t\in\mathbb{R}^{+}$.
It is important to clarify the physical interpretation of (9). In a general time-varying velocity field, the position of a particle is found by integrating the velocity field over its trajectory. Our definition provides a simplification of this picture, where the total displacement of the field at location $\mathbf{s}$ and time $t$ is determined directly by the velocity vector at that final space-time coordinate. This makes the model analytically tractable while still capturing the essence of a locally varying advection.
The velocity field $\mathbf{v}(\mathbf{s},t)$ is an interpretable, infinite-dimensional parameter that connects our model to established concepts in fluid mechanics (kundu2015fluid). The Evolving FF is a flexible class of models that includes several well-known flows as special cases. For instance, if the velocity is constant in time, $\frac{\partial}{\partial t}\mathbf{v}(\mathbf{s},t)=\mathbf{0}$, this defines a steady flow. Examples of such steady flows include a rigid body rotation $\mathbf{v}(\mathbf{s})=(-ks_{2},ks_{1})$, a stagnation point $\mathbf{v}(\mathbf{s})=(ks_{1},-ks_{2})$, and a vortex $\mathbf{v}(\mathbf{s})=(-ks_{2}/s,ks_{1}/s)$ for $s=\sqrt{s_{1}^{2}+s_{2}^{2}}$. Other important special cases are defined by physical constraints, such as incompressible flow ($\nabla\cdot\mathbf{v}=\mathbf{0}$) or irrotational flow ($\nabla\times\mathbf{v}=\mathbf{0}$). Alternatively, a simpler non-stationary model assumes the velocity is constant in space, $\mathbf{v}(\mathbf{s},t)=\mathbf{v}(t)$.
Hence, model (9) is highly relevant, and determining its mathematical properties is intrinsecally linked to the unobserved velocity field $\mathbf{v}(\mathbf{s},t)$.
Proposition 3.2.
Assume that $Z(\mathbf{s},t)$ is an evolving frozen field as specified by Definition 3.2. For fixed $(\mathbf{s},t)\in\mathbb{R}^{2}\times\mathbb{R}^{+}$, we have
$\mathrm{cov}(Z(\mathbf{s},t),Z(\mathbf{s}+\mathbf{h},t+\tau))=$
$c_{XX}\left(\mathbf{h}-\mathbf{v}(\mathbf{s},t)\tau-\left\{{\mathbf{D}}_{\mathbf{s}}(\mathbf{s},t)\mathbf{h}+D_{t}(\mathbf{s},t)\tau+O(\|\mathbf{h}\|^{2})+O(\tau^{2})\right\}(t+\tau)\right),$ (10)
with ${\mathbf{D}}_{\mathbf{s}}(\mathbf{s},t)=\begin{pmatrix}\frac{\partial\text{v}_{1}(\mathbf{s},t)}{\partial s_{1}}&\frac{\partial\text{v}_{1}(\mathbf{s},t)}{\partial s_{2}}\\ \frac{\partial\text{v}_{2}(\mathbf{s},t)}{\partial s_{1}}&\frac{\partial\text{v}_{2}(\mathbf{s},t)}{\partial s_{2}}\end{pmatrix}$ and $D_{t}(\mathbf{s},t)=\left(\frac{\partial\text{v}_{1}(\mathbf{s},t)}{\partial t},\frac{\partial\text{v}_{2}(\mathbf{s},t)}{\partial t}\right)^{T}$. Moreover, the propagation path of $Z(\mathbf{s},t)$ is
$\left((\mathbf{I}-{\mathbf{D}}_{\mathbf{s}}(\mathbf{s},t)(t+\tau))^{-1}\left\{\mathbf{v}(\mathbf{s},t)+D_{t}(\mathbf{s},t)(t+\tau)\right\}\tau,\tau\right),$ (11)
where $\mathbf{I}$ is the identity matrix of dimension $2$.
Proof.
First, we apply local expansions on $\mathbf{v}(\mathbf{s},t)$. For fixed $(\mathbf{s},t)\in\mathbb{R}^{2}\times\mathbb{R}^{+}$, we express the velocity field locally to be:
$\mathbf{v}(\mathbf{s}+\mathbf{h},t+\tau)=\mathbf{v}(\mathbf{s},t)+{\mathbf{D}}_{\mathbf{s}}(\mathbf{s},t)\mathbf{h}+D_{t}\tau+O(\|\mathbf{h}\|^{2})+O(\tau^{2}).$ (12)
Then, we plug the expansion (12) into the covariance of the evolving frozen field, i.e.
$\mathrm{cov}(Z(\mathbf{s},t),Z(\mathbf{s}+\mathbf{h},t+\tau))=$
$\mathrm{cov}\left\{X(\mathbf{s}-\mathbf{v}(\mathbf{s},t)t),X(\mathbf{s}+\mathbf{h}-\mathbf{v}(\mathbf{s}+\mathbf{h},t+\tau)(t+\tau))\right\}=$
$c_{XX}\left(\mathbf{h}-\left(\mathbf{v}(\mathbf{s}+\mathbf{h},t+\tau)(t+\tau)-\mathbf{v}(\mathbf{s},t)t\right)\right)=$
$c_{XX}\left(\mathbf{h}-\left\{\mathbf{v}(\mathbf{s},t)+{\mathbf{D}}_{\mathbf{s}}(\mathbf{s},t)\mathbf{h}+D_{t}(\mathbf{s},t)\tau+O(\|\mathbf{h}\|^{2})+O(\tau^{2})\right\}(t+\tau)+\mathbf{v}(\mathbf{s},t)t\right)=$
$c_{XX}\left(\mathbf{h}-\mathbf{v}(\mathbf{s},t)\tau-\left\{{\mathbf{D}}_{\mathbf{s}}(\mathbf{s},t)\mathbf{h}+D_{t}(\mathbf{s},t)\tau+O(\|\mathbf{h}\|^{2})+O(\tau^{2})\right\}(t+\tau)\right).$ (13)
Finally, setting the argument of (3.1) to zero yields the propagation path. ∎
Notice that the resulting covariance (3.2) and propagation path (11) are local analytical approximations, valid for small spatial and temporal lags.
In summary, with the models presented in this section, we avoid having perfect correlation along $(\mathbf{v}\tau,\tau)$ in two ways: either by having random rotations and scaling of the constant velocity using transformations (5), or by “perturbing” the propagation path as in (11) by adding higher-order terms.
3.2 Simulated Examples of a Cyclone
To illustrate the behavior of these velocity-based generalizations, we now present simulation examples for both the Distributed and Evolving FF models. We choose to simulate the evolution of a tropical cyclone (dvorak1975tropical), which has attracted a lot of interest in the climatology community for parametric models used in estimation (see for instance olander2007, knaff2010, pradhan2018, maskey2020). Our aim is to show that our proposed models can simulate complex dynamics that go well beyond those allowed by the FF model, and that are of interest in the climate science community.
Before discussing the simulated examples, we illustrate how to simulate from the two models introduced in Section 3.1. First, the Distributed FF consists in a convex combination of $n_{v}$ realizations of FFs, each with their specified velocity and wight. Thus, to simulate a Distributed FF one just simulates each FF $n_{v}$ times with Algorithm 1. Moreover, the simulation of an Evolving FF can also be carried out with a modification of Algorithm 1. Specifically, when the space random field $X(\mathbf{s})$ is moved pointwisely in time, instead of shifting it according to a constant velocity, it is moved according to a velocity field $\mathbf{v}(\mathbf{s},t)$. Being it a continuous function of space and time, $\mathbf{v}(\mathbf{s},t)$ is discretized over the space and time grids. For practical implementation, we suggest discretizing $\mathbf{v}(\mathbf{s},t)$ on the same space-time grid as the simulation. When it comes to establishing how large grid $\mathcal{G}^{*}$, should be, one can adopt $\text{v}_{\text{max}}=\max\{\lVert\text{v}_{1}\rVert_{\infty},\lVert\text{v}_{2}\rVert_{\infty}\}$ for varying velocity $\mathbf{v}(\mathbf{s},t)=(\text{v}_{1}(\mathbf{s},t),\text{v}_{2}(\mathbf{s},t))^{T}$, where we determine $\lVert\text{v}_{i}\rVert_{\infty}=\inf\{V\geq 0:|\text{v}_{i}|\leq V\text{ for a.e. }(\mathbf{s},t)\}$, $i=1,2$. We include both algorithms in the supplementary materials.
We now illustrate the Distributed FF concept with a simulation where the component velocities are resampled at each time step, allowing the overall pattern to evolve. In particular, for a given time $t\in\{0,\dots,T\}$, we are interested in combining rotation and translation of the random field. We first draw a set of $n_{\theta}$ random angles,
$\theta_{1}^{t},\dots,\theta_{n_{\theta}}^{t}\stackrel{{\scriptstyle iid}}{{\sim}}f_{R}(\theta;\boldsymbol{\alpha}_{R}),$ (14)
and a single random velocity vector,
$\mathbf{v}^{t}\sim f_{T}(\mathbf{v};\boldsymbol{\alpha}_{T}).$ (15)
Equation (14) denotes the $n_{\theta}$ angles that define the random rotations of the field, where $f_{R}$ is a probability density function (PDF). Equation (15) describes the velocity $\mathbf{v}^{t}$ with which the entire field is advected at time $t$.
For a fixed $t$, we obtain the distributed frozen field as a weighted average of multiple rotated and translated fields:
$Z(\mathbf{s},t)$ $=\sum_{i=1}^{n_{\theta}}p(\theta^{t}_{i})X^{\theta^{t}_{i}}_{S}(\mathbf{s}-\mathbf{v}^{t}t).$
Here, $X^{\theta^{t}_{i}}_{S}$ denotes the base random field $X_{S}$ after being rotated by the angle $\theta^{t}_{i}$ around a fixed center of rotation, $C$. The term $(\mathbf{s}-\mathbf{v}^{t}t)$ indicates that this rotated field is evaluated at a coordinate shifted by the velocity vector, effectively modeling advection. The weights $p(\theta^{t}_{i})$ are constructed to form a convex combination, ensuring that $p(\theta^{t}_{i})\geq 0$ and $\sum_{i=1}^{n_{\theta}}p(\theta^{t}_{i})=1$. They are derived from the PDF, $f_{R}$, evaluated at each sampled angle. Specifically, the weight for each rotation is calculated by normalizing the PDF value at that angle:
$p(\theta^{t}_{i})=\frac{f_{R}(\theta^{t}_{i};\boldsymbol{\alpha}_{R})}{\sum_{j=1}^{n_{\theta}}f_{R}(\theta_{j}^{t};\boldsymbol{\alpha}_{R})}.$
This approach ensures that rotations corresponding to angles with higher probability density are given more weight in the final mixture.
To simulate the example in Figure 1, we set $n_{\theta}=10$. For each $t\in\{0,\dots,7\}$, the angles are drawn from a Wrapped Normal distribution, $\theta^{t}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{WN}(0,5)$, and the velocity is drawn from a multivariate Normal distribution, $\mathbf{v}^{t}\sim\mathcal{N}(\boldsymbol{\mu},\mathbf{I})$, with mean $\boldsymbol{\mu}=(5,0)^{T}$ and the identity as covaraince matrix. The rotation is performed around a fixed center $C=(50,65)$. The x marks in the plot indicate the position of this rotation center relative to the advected field, given by $C^{t}=C+\mathbf{v}^{t}t$. For this simulation, the initial field at $t=0$, $Z(\boldsymbol{s},0)=X(\boldsymbol{s})$, is a fixed, deterministic image rather than a realization of a random field. We deliberately choose a simple shape, a faded circle on a constant background, to provide a clear visual benchmark. This approach allows the model’s intrinsic dynamics to be observed directly and clearly.
Figure 1: Distributed FF simulation of a cyclone.
It could be of interest to simulate how the cyclone evolves locally, namely how a single portion of it travels and gets deformed along time. We simulate such trajectory with an Evolving FF model. In particular, Figure 2 shows an example in which velocity field $\mathbf{v}(\mathbf{s},t)$ is such that
$\mathbf{v}(\mathbf{s},t)=\left(\frac{R(\mathbf{s})\cdot\text{cos}(t)}{t+1},\frac{R(\mathbf{s})\cdot\text{sin}(t)}{t+1}\right)^{T},$ (16)
with
[…]
Velocity field (16) describes a spiral whose radius $R(\mathbf{s})$ changes at each spatial location. Hence, the magnitude of the point-wise elements of the velocity field depends on both the location and the time, while their direction only on the latter. In the top and bottom rows of Figure 2 we show the traveling field and the corresponding velocity fields changing over time, respectively. Notice that we used the same initial image as in Figure 1, and that the velocity fields have been re-scaled for representation purposes. As we would expect in the observed temporal evolution of a tropical cyclone, the points closest to the center, i.e. the origin, get less distorted than those further away from it.
Figure 2: Evolving FF simulation of a cyclone (top row) and the corresponding velocity filed $\mathbf{v}(\mathbf{s},t)$ at every time point (bottom row).
3.3 Comparison with Numerical Solvers
In this section we show how our proposed methods can simulate solutions of some classes of Partial Differential Equations (PDEs). A comparison with numerical PDE solvers is particularly appropriate, as both our method and these solvers are fundamentally constructed upon a specified advection velocity. This shared foundation provides a direct and rigorous basis for validation. We consider simulations on a spatial grid $\mathcal{G}=[0,100]\times[0,100]$ and time grid $t\in\{0,1,2,3\}$. We use package ReacTran (reactran) to compute the solution of the PDEs. Moreover, to compare the simulated space-time random field $Z(\mathbf{s},t)$ and the solution of the numerical solver $Z_{\text{PDE}}(\mathbf{s},t)$, we use Root Mean Square Error (RMSE)
$\text{RMSE}(t)=\sqrt{\frac{1}{n^{2}}\sum_{\mathbf{s}\in\mathcal{G}}\left(Z(\mathbf{s},t)-Z_{\text{PDE}}(\mathbf{s},t)\right)^{2}},\quad t\in\{0,1,2,3\}.$
As mentioned in Section 2.1, the FF corresponds to the solution of the pure transport PDE (2). Thus, the most natural first comparison we make is between our simulated FF and the solution of a transport equation. We simulate a FF with $\mathbf{v}=(20,10)^{\top}$, and we solve (2) with a no-flux boundary condition. The results are shown in Figure 3. We notice that our proposed simulation method closely resembles the numerical solution of the PDE. However, the latter shows numerical diffusion, so that the concentration of the image gets slightly dissipated with time (see for instance LeVeque_2002). This dissipation is quantified in the RMSE plot. the error is zero at $t=0$, as both methods are initialized with the identical field, but grows over time due to the numerical diffusion in the PDE solver. By construction, our suggested method described in Algorithm 1 does not suffer from this issue.
Figure 3: On the left, solution of the transport PDE (top row) and the corresponding simulated Frozen Field (bottom row). On the right, the RMSE between two images at each time point.
We now turn to the comparison of a Distributed FF with the solution of a transport-diffusion PDE. We show how the superposition of multiple FFs, each with a distinct advection velocity, gives rise to both a mean advective transport and an effective diffusive term. We begin by considering that the field is a superposition of multiple components, each transported by a distinct advection velocity. Let this set of component velocities be denoted by $\{\mathbf{v}_{i}\}_{i=1}^{n_{v}}$, with each velocity occurring with a corresponding probability $\{p_{i}\}_{i=1}^{n_{v}}$. The mean velocity of the system, $\overline{\mathbf{v}}$, is the weighted average of these component velocities, namely $\overline{\mathbf{v}}=\sum_{i=1}^{n_{v}}p_{i}\mathbf{v}_{i}$. The velocity deviations from this mean are then defined as $\Delta\mathbf{v}_{i}=\mathbf{v}_{i}-\overline{\mathbf{v}}$. Conceptually, each component velocities $\mathbf{v}_{i}$ can be thought of as arising from a physical transformation $\mathbf{M}_{i}$ acting on a base velocity $\mathbf{v}$. Similarly to christakos2017spatiotemporal, treating the velocity as the sum of a mean and the fluctuations around is an advantageous framework, as we shall see next.
It can be shown (see supplementary materials) that for fixed $(\mathbf{s},t)$, the governing equation for the DFF field $Z(\mathbf{s},t)$ is:
$\frac{\partial Z(\mathbf{s},t)}{\partial t}+\overline{\mathbf{v}}\cdot\nabla Z(\mathbf{s},t)=-\nabla\cdot\left(\sum_{i=1}^{n_{v}}p_{i}\,Z_{i}(\mathbf{s},t)\Delta\mathbf{v}_{i}\right).$ (17)
The left-hand side of (17) represents the pure advection of the field $Z$ at the mean velocity $\overline{\mathbf{v}}$. The term on the right is the divergence of the “fluctuating flux”, which arises from the velocity deviations. The equation in its current form is not closed, as the flux depends on the individual components $Z_{i}$. To create a macroscopic equation for $Z$, we must impose further assumptions.
We adopt the widely used gradient-diffusion hypothesis, which assumes that the fluctuating flux is proportional to the negative gradient of the mean field. In other words, this implies that the random motion of particles will naturally cause them to move from a more crowded region to a less crowded one. In our model, this entails
$\sum_{i=1}^{n_{v}}p_{i}\,Z_{i}(\mathbf{s},t)\Delta\mathbf{v}_{i}=-\mathbf{C}\nabla Z(\mathbf{s},t),$ (18)
where $\mathbf{C}$ is the diffusion tensor. Substituting (18) into (17), we arrive at the advection-diffusion equation:
$\frac{\partial Z(\mathbf{s},t)}{\partial t}+\overline{\mathbf{v}}\cdot\nabla Z(\mathbf{s},t)=\nabla\cdot(\mathbf{C}\nabla Z(\mathbf{s},t)).$
This derivation provides a formal link between our Distributed FF model and the advection-diffusion PDE.
The identification of the diffusion tensor $\mathbf{C}$ with the covariance of the velocity distribution (Taylor1954) has a physical interpretation. Specifically, diffusion is a macroscopic process of spatial spreading. In the Distributed FF model, the microscopic cause of this spreading is the fact that different additive components of the field are being transported at different velocities. The velocity covariance matrix, namely
$\mathbf{C}_{\mathbf{v}}=\sum_{i=1}^{n_{v}}p_{i}\,\Delta\mathbf{v}_{i}\Delta\mathbf{v}_{i}^{\top}=\sum_{i=1}^{n_{v}}p_{i}\,(\mathbf{v}_{i}-\overline{\mathbf{v}})(\mathbf{v}_{i}-\overline{\mathbf{v}})^{\top},$
is the precise mathematical tool that quantifies the magnitude and orientation of this velocity spread. Therefore, the diffusion tensor $\mathbf{C}$ in the macroscopic PDE is the direct physical manifestation of the velocity covariance, and we can assume that $\mathbf{C}$ is proportional to $\mathbf{C}_{\mathbf{v}}$. This principle allows us to construct a Distributed FF that mimics a desired advection-diffusion process by defining a velocity distribution with the appropriate mean and covariance.
To validate our theoretical framework, we compare our Distributed FF model against a numerical solution of the advection-diffusion equation. The PDE was solved on the same spatial and time grids as in the previous FF example, with a velocity of $\overline{\mathbf{v}}=(20,10)^{T}$ and a diffusion tensor of $\mathbf{C}=\begin{pmatrix}5&0\\ 0&1.5\end{pmatrix}$. The Distributed FF model was constructed by drawing $\{\mathbf{v}_{i}\}_{i=1}^{n_{v}}$ from a multivariate normal distribution whose parameters were set to match the PDE, i.e. $\mathbf{v}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(\overline{\mathbf{v}},\mathbf{C})$. The weights used are $p_{i}=1/n_{v}$, for each $i=1,\dots,n_{v}$. In Figure 4 we compare single Distributed FF realization with $n_{v}=50$ and the PDE numerical solution, showing qualitative agreement between the two. To quantify the model’s convergence and performance, we carry out a simulation study of $500$ iterations using different numbers of components, $n_{v}=10,50,200$. The right panel of Figure 4 quantifies the model’s performance against the PDE solution. The results clearly indicate that as $n_{v}$ increases, both the median RMSE and its variance decrease across all time points, demonstrating that the Distributed FF model converges to the true solution as more components are included. The shape of the RMSE curves also provides insight into the process. As established in the pure transport case, the error is zero at $t=0$ because both methods are initialized with the identical field. The RMSE then rises as the approximation error accumulates, particularly during the initial phase where the concentration gradients are steepest. Subsequently, the curve plateaus because the diffusion process itself smooths the field over time. As the gradients lessen, the rate at which new error is generated decreases, causing the total error to stabilize.
Figure 4: On the left, solution of the transport-diffusion PDE (top row) and a simulated Distributed Frozen Field with $n_{v}=50$. (bottom row). On the right, the boxplots of RMSE between two images at each time point across 500500 iterations and for $n_{v}=10,50,200$.
4 Generalization via Relaxation of the Spectrum Shape
In this section, we consider a different type of generalization of the FF model, namely by relaxing the shape of its spectrum (4). We first propose a new process by defining its spectrum. Then, we analyze the role of the parameters in such definition with an example.
4.1 Framework
For a fixed wavenumber $\mathbf{k}$ and velocity $\mathbf{v}$, the spectrum of a FF (4) consists of the spectrum $S_{XX}(\mathbf{k})$ of the spatial random field multiplied by a Dirac delta centered in $\omega=-\mathbf{k}^{T}\mathbf{v}$. Such separable structure characterizing $S_{ZZ}(\mathbf{k},\omega)$ makes it particularly convenient for us to modify the part of the spectrum that depends on velocity $\mathbf{v}$. For instance, replacing the Dirac deltas with a “milder” singularity would ensure that the spectrum includes more frequencies neighbouring $\omega=-\mathbf{k}^{T}\mathbf{v}$. This introduces temporal decorrelation into the model, which gains the flexibility to represent more complex and realistic phenomena where features naturally change, dissipate, or evolve over time.
The following proposition provides the theoretical formalization for this intuition, proving that for a broad class of stationary fields, perfect temporal correlation is a unique property of the FF model.
Proposition 4.1.
Let $Z(\mathbf{s},t)$ be a stationary space-time random field with a valid spectrum $S_{ZZ}(\mathbf{k},\omega)$ and covariance function $c_{ZZ}(\mathbf{h},\tau)$. Then, for any $\tau>0$, the correlation along the path $(\mathbf{v}\tau,\tau)$ satisfies
$\rho(\mathbf{v}\tau,\tau)=\frac{|c_{ZZ}(\mathbf{v}\tau,\tau)|}{c_{ZZ}(0,0)}<1,$
unless $Z(\mathbf{s},t)$ is a frozen field with velocity $\mathbf{v}$.
Proof.
By definition, the covariance $c_{ZZ}(\mathbf{h},\tau)$ is the inverse Fourier transform of the spectrum $S_{ZZ}(\mathbf{k},\omega)$:
$c_{ZZ}(\mathbf{h},\tau)=\iiint S_{ZZ}(\mathbf{k},\omega)\,e^{2\pi i(\mathbf{k}^{\top}\mathbf{h}+\omega\tau)}\,\mathrm{d}\omega\,\mathrm{d}^{2}\mathbf{k}.$
Evaluating this along the path $(\mathbf{v}\tau,\tau)$ and at the origin gives:
$c_{ZZ}(\mathbf{v}\tau,\tau)$ $=\iiint S_{ZZ}(\mathbf{k},\omega)\,e^{2\pi i\tau(\mathbf{k}^{\top}\mathbf{v}+\omega)}\,\mathrm{d}\omega\,\mathrm{d}^{2}\mathbf{k},$
$c_{ZZ}(0,0)$ $=\iiint S_{ZZ}(\mathbf{k},\omega)\,\mathrm{d}\omega\,\mathrm{d}^{2}\mathbf{k}.$
By the triangle inequality, we have:
$\left|\iiint S_{ZZ}(\mathbf{k},\omega)\,e^{2\pi i\tau(\mathbf{k}^{\top}\mathbf{v}+\omega)}\,\mathrm{d}\omega\,\mathrm{d}^{2}\mathbf{k}\right|\leq\iiint S_{ZZ}(\mathbf{k},\omega)\,\left|e^{2\pi i\tau(\mathbf{k}^{\top}\mathbf{v}+\omega)}\right|\,\mathrm{d}\omega\,\mathrm{d}^{2}\mathbf{k}.$
Since $|e^{i\theta}|=1$ for any real $\theta\in\mathbb{R}$, the inequality simplifies to $|c_{ZZ}(\mathbf{v}\tau,\tau)|\leq c_{ZZ}(0,0)$. Equality can only hold if the argument of the integral, $S_{ZZ}(\mathbf{k},\omega)e^{2\pi i\tau(\mathbf{k}^{\top}\mathbf{v}+\omega)}$, has a constant phase for all $(\mathbf{k},\omega)$ where the spectrum $S_{ZZ}$ is non-zero. Given that $S_{ZZ}(\mathbf{k},\omega)$ is a real and non-negative function, it does not affect the phase. Therefore, the phase of the complex exponential term $e^{2\pi i\tau(\mathbf{k}^{\top}\mathbf{v}+\omega)}$ must be constant, which implies that its exponent, $2\pi\tau(\mathbf{k}^{\top}\mathbf{v}+\omega)$, must be constant. This condition must hold for all $(\mathbf{k},\omega)$ with non-zero spectral energy. If we consider the point $(\mathbf{k}=\mathbf{0},\omega=0)$, which must also satisfy this relation, we find that $2\pi\tau(0+0)=0$. Thus, with the constant being zero, the condition for equality becomes $\mathbf{k}^{\top}\mathbf{v}+\omega=0$ on the set of wavenumber-frequency $\{(\mathbf{k},\omega)\in\mathbb{R}^{2}\times\mathbb{R}:S_{ZZ}(\mathbf{k},\omega)>0\}$, implying $\omega=-\mathbf{k}^{\top}\mathbf{v}$ in the same set. This means the spectrum $S_{ZZ}$ must be entirely supported on the hyperplane defined by $\omega=-\mathbf{k}^{\top}\mathbf{v}$, which is precisely the spectral definition of a FF.
Proposition 4.1 formally establishes that the perfect temporal correlation of the FF model is an idealized case existing only under a strict spectral condition. The proof guarantees that any stationary model with advection that deviates even slightly from the FF’s rigid structure will inherently exhibit temporal decorrelation along a propagation path. Among the models in this class, we build one where the Dirac delta is substituted with a decaying function that distributes power to a broader range of frequencies around the central frequency $\omega=-\mathbf{k}^{\top}\mathbf{v}$. This choice, inspired by seasonally persistent processes (andvel1986long; gray1989generalized) and models of geophysical phenomena (stein1999interpolation), has the advantage of preserving the interpretability of the FF while being more physically plausible. The following definition introduces a unified model.
Definition 4.1 (Damped FF).
A Damped Frozen Field is a zero-mean Gaussian process characterized by a persistence parameter $\alpha>0$ and a damping parameter $\beta\geq 0$, whose spectrum is defined as
$S_{ZZ}(\mathbf{k},\omega)=S_{XX}(\mathbf{k})\cdot B(\mathbf{k},\omega;\mathbf{v},\beta)^{-\alpha},$ (19)
where
$B(\mathbf{k},\omega;\mathbf{v},\beta)=(\omega+\mathbf{k}^{\top}\mathbf{v})^{2}+(\beta\omega)^{2}.$
This model captures the interplay between advection and physical decorrelation through its two key parameters. The persistence parameter, $\alpha$, controls the strength of the process’s memory along the advective path. The damping parameter, $\beta$, prevents the singularity of a FF, and governs the geometric shape of the spectral decay around the advection line. Notice that persistence parameter $\alpha$ must be positive. This ensures that the model describes a system where energy decays at higher frequencies, which is a necessary condition for a well-behaved physical process. Furthermore, for the total variance of the process to be finite, i.e. $c_{ZZ}(0,0)=\iiint S_{ZZ}(\mathbf{k},\omega)\,\mathrm{d}\omega\,\mathrm{d}^{2}\mathbf{k}<\infty$, the spectrum must be integrable. This typically imposes a stronger lower bound on $\alpha$, depending on the behavior of the spatial spectrum $S_{XX}(\mathbf{k})$ as $\mathbf{k}\to 0$. We will discuss further the roles of parameters $\alpha$ and $\beta$ in Section 4.2.
Proposition 4.2.
The spectrum $S_{ZZ}(\mathbf{k},\omega)$ specified in Definition 4.1 corresponds to a valid, positive definite covariance function.
Proof.
This follows from Bochner’s theorem (priestley1981spectral). The spectrum $S_{ZZ}(\mathbf{k},\omega)$ is non-negative, as it is the product of the non-negative spatial spectrum $S_{XX}(\mathbf{k})$ and the strictly positive damping factor $B(\mathbf{k},\omega;\mathbf{v},\beta)^{-\alpha}$. Since the Fourier transform establishes a one-to-one correspondence between non-negative spectra and positive definite covariance functions, $c_{ZZ}(\mathbf{h},\tau)$ is guaranteed to be valid. ∎
Finally, the simulation of the Damped FF is done directly with Algorithm 2, as this method can generate a realization from any valid, specified spectrum. It suffices to substitute the Damped FF spectrum from Equation (19) into the algorithm, bearing in mind the practical need to define the spectrum on a discrete frequency grid. We include the algorithm in the supplementary materials.
4.2 Role of the Parameters
The spectrum in Equation (19) is shaped by the term $B(\mathbf{k},\omega;\mathbf{v},\beta)^{-\alpha}$. The power-law form of this term is characteristic of a long-memory process, but with the critical distinction that its behavior is centered on the advection line $\omega=-\mathbf{k}^{\top}\mathbf{v}$ rather than the origin. This formulation defines a space-time long-memory process with advection, whose characteristics are non-trivial due to the non-separable nature of $B(\mathbf{k},\omega;\mathbf{v},\beta)$. This term acts as a filter that intrinsically links a wave’s temporal evolution to its spatial wavenumber, which is a physically realistic assumption. The filter reshapes the initial spatial spectrum by amplifying the signal where $B(\mathbf{k},\omega;\mathbf{v},\beta)<1$, and attenuating it where $B(\mathbf{k},\omega;\mathbf{v},\beta)>1$. The geometry of this amplification region is primarily controlled by the damping parameter $\beta$, while the persistence parameter $\alpha$ dictates the intensity of the effect.
To illustrate this, we set $\mathbf{v}=(5,0)^{\top}$ and, by considering only the plane where $k_{2}=0$, we visualize the behavior of $B(\mathbf{k},\omega;\mathbf{v},\beta)$ over the domain $k_{1}\in[-\pi/4,\pi/4]$ and $\omega\in[-\pi,\pi]$. Figure 5 illustrates the interplay between damping and persistence parameters, and the relative effects on $\log_{10}\left(B(\mathbf{k},\omega;\mathbf{v},\beta)^{-\alpha}\right)$. As mentioned previously, $\beta$ governs the geometric shape of the persistent region in the frequency domain. When $\beta=0$, the filter creates an unbounded corridor of perfect persistence along the advection line $\omega=-\mathbf{k}^{\top}\mathbf{v}$. However, when $\beta>0$, the $(\beta\omega)^{2}$ term effectively dampens low temporal frequencies. This closes the corridor into a bounded, elliptical, region, meaning only a specific band of frequencies is allowed to persist over time. In turn, the parameter $\alpha$ controls the degree of this persistence by dictating the filter’s selectivity. A higher $\alpha$ makes the filter more selective, as it amplifies frequencies inside the persistent region defined by $\beta$, and attenuates those outside. This makes the process resemble a pure advective field. Thus, we conclude that the persistence parameter controls the long-range temporal memory of the process.
In summary, the damping and persistence parameters offer a flexible framework to model a rich variety of space-time processes by capturing the interplay between advection and decay of the correlation.
Figure 5: Heatmap of the logarithm of $B(\mathbf{k},\omega;\mathbf{v},\beta)^{-\alpha}$ for different choices of $\alpha$ (rows) and $\beta$ (columns). Level curves are shown in a gradient of pinks, where lightest correspond to highest values. The dashed black line correspond to $\omega=-\mathbf{k}^{\top}\mathbf{v}$.
5 Conclusions
In this work, we have developed flexible and physically interpretable models for spatio-temporal phenomena governed by advection. Recognizing the limitations of the classical FF model, especially its assumption of perfect correlation along the advection path, we introduced two distinct generalization strategies: one based on enriching the velocity structure, and the other on relaxing the spectral shape.
Our first strategy led to the development of the Distributed FF and the Evolving FF models. The Distributed FF, by representing the process as a superposition of fields with different advective velocities, provides a natural framework for modeling phenomena driven by multiple transport influences. We demonstrated that this model is not only capable of producing complex dynamics, such as those of a tropical cyclone, but also that it provides a direct stochastic analogue to the advection-diffusion PDE. The Evolving FF, in contrast, allows the velocity to vary locally in space and time, offering a powerful tool for simulating processes within complex, non-uniform flow fields. For both models, we derived key properties, including their propagation paths, and provided efficient simulation algorithms based on modifications of the FF simulation.
Our second strategy directly addressed the spectral properties of the advective process. We introduced the Damped FF model, which replaces the singular Dirac delta in the FF spectrum with a flexible, continuous filter. This filter is governed by a persistence parameter, $\alpha$, and a damping parameter, $\beta$, which together control the rate of decorrelation and the geometric shape of the spectral power around the advection line $\omega=-\mathbf{k}^{\top}\mathbf{v}$. We proved that this construction yields a valid covariance function and explored how the interplay between $\alpha$ and $\beta$ allows for the modeling of a rich class of processes that bridge the gap between pure advection and rapid decay.
The frameworks established in this paper open several promising directions for future investigation. First, while our work has focused on model definition and simulation, an avenue for future research is the development of robust statistical methods for fitting these models to observational data. This includes estimating the velocity distribution for a Distributed FF, the latent velocity field for an Evolving FF, or the spectral parameters for a Damped FF. Techniques ranging from Whittle likelihood estimation (guillaumin2019debiased) to hierarchical Bayesian modeling (see for instance rue2007 and chagneau2011) could be explored for this purpose. Once estimation procedures are developed, a crucial line of inquiry will be the direct comparison of these models against competing approaches in the literature. Such a comparison should focus on predictive performance and forecasting capabilities, which are critical for real-world applications in fields like climatology, atmospheric science, and oceanography. Second, many environmental systems involve multiple, co-evolving variables, like temperature and pressure, and pollutants and wind velocity. Extending our models to a multivariate setting, where several fields are jointly advected and interact, is a critical next step. The multivariate FF model of salvana2022jasa provides a foundational starting point for developing multivariate versions of the Distributed, Evolving, and Damped FF models. Finally, the Damped FF model introduces a form of long memory centered on the advection path. A deeper theoretical investigation could formally connect this model to the theories of fractional Brownian motion (e.g., flandrin1989; qian1998two; stein2002fast; lilly2017fractional; didier2018domain) and spatial long-memory processes (robinson2020spatial). For instance, it would be valuable to understand the interaction between the long-memory characteristics of the initial spatial field $X_{S}(\mathbf{s})$ and the long-memory properties induced by the advective spectral filter.
6 Declaration of Interest
None.
7 Acknowledgments
Sofia Olhede would like to thank the European Research Council under Grant CoG 2015-682172NETS, within the Seventh European Union Framework Program.
Appendix A Algorithms
This section outlines the simulation algorithms for the models presented in the main paper. Algorithm 3 shows how to simulate a Distributed Frozen Field as a convex combination of $n_{v}$ simulated Frozen Fields; each component $i=1,\dots,n_{v}$ is advected with a unique velocity $\mathbf{v}_{i}=\mathbf{M}_{i}\mathbf{v}$ and assigned a weight $p_{i}=p(\mathbf{M}_{i}\mathbf{v})$. For the Evolving Frozen Field, Algorithm 4 adapts the simulation procedure of a Frozen Field by substituting the constant advection velocity with a vector field $\mathbf{v}(\mathbf{s},t)$ that varies for each spatial location $\mathbf{s}$ and time point $t$. Finally, Algorithm 5 presents the spectral simulation of a Damped Frozen Field, which is generated from a spectrum constructed using the spatial field’s spectrum $S_{XX}$, and a filter whose shape is determined by a persistence parameter $\alpha>0$ and a damping parameter $\beta\geq 0$.
Algorithm 3 Simulation of a Distributed Frozen Field via Grid Shifting
1:Spatial grid dimension $n$, number of time steps $T$, base velocity $\mathbf{v}$, spatial covariance function $c_{XX}$.
2:Number of components $n_{v}$, transformation matrices $\{\mathbf{M}_{1},\dots,\mathbf{M}_{n_{v}}\}$, weights $\{p_{1},\dots,p_{n_{v}}\}$.
3:Compute transformed velocities $\mathbf{v}_{i}\leftarrow\mathbf{M}_{i}\mathbf{v}$ for $i=1,\dots,n_{v}$.
4:Compute max velocity component $v_{\text{max}}\leftarrow\max_{i\in\{1,\dots,n_{v}\},j\in\{1,2\}}\{|(\mathbf{v}_{i})_{j}|\}$.
5:Compute extended grid dimension $N\leftarrow n+Tv_{\text{max}}$.
6:Define extended grid $\mathcal{G}^{*}\leftarrow[0,N]\times[0,N]$.
7:Generate Gaussian random field $X_{S}^{\text{ext}}$ on $\mathcal{G}^{*}$ using $c_{XX}$.
8:Define simulation grid $\mathcal{G}\leftarrow[0,n]\times[0,n]$.
9:Initialize space-time field $Z\in\mathbb{R}^{(n+1)\times(n+1)\times(T+1)}$.
10:for each spatial location $\mathbf{s}$ in $\mathcal{G}$ do
11:  $Z(\mathbf{s},0)\leftarrow X_{S}^{\text{ext}}(\mathbf{s})$.
12:  for $t=1$ to $T$ do
13:   $\text{value}\leftarrow 0$.
14:   for $i=1$ to $n_{v}$ do
15:     $\text{value}\leftarrow\text{value}+p_{i}\cdot X_{S}^{\text{ext}}(\mathbf{s}-\mathbf{v}_{i}t)$.
16:   end for
17:   $Z(\mathbf{s},t)\leftarrow\text{value}$.
18:  end for
19:end for
20:return $Z$.
Algorithm 4 Simulation of an Evolving Frozen Field via Grid Shifting
1:Spatial grid dimension $n$, number of time steps $T$, velocity field $\mathbf{v}(\mathbf{s},t)$, spatial covariance function $c_{XX}$.
2:Define simulation grid $\mathcal{G}\leftarrow[0,n]\times[0,n]$.
3:Compute max velocity component $v_{\text{max}}\leftarrow\max_{\mathbf{s}\in\mathcal{G},t\in\{1,\dots,T\},j\in\{1,2\}}\{|v_{j}(\mathbf{s},t)|\}$.
4:Compute extended grid dimension $N\leftarrow n+Tv_{\text{max}}$.
5:Define extended grid $\mathcal{G}^{*}\leftarrow[0,N]\times[0,N]$.
6:Generate Gaussian random field $X_{S}^{\text{ext}}$ on $\mathcal{G}^{*}$ using $c_{XX}$.
7:Initialize space-time field $Z\in\mathbb{R}^{(n+1)\times(n+1)\times(T+1)}$.
8:for each spatial location $\mathbf{s}$ in $\mathcal{G}$ do
9:  $Z(\mathbf{s},0)\leftarrow X_{S}^{\text{ext}}(\mathbf{s})$.
10:  for $t=1$ to $T$ do
11:   $\mathbf{v}_{st}\leftarrow\mathbf{v}(\mathbf{s},t)$.
12:   $Z(\mathbf{s},t)\leftarrow X_{S}^{\text{ext}}(\mathbf{s}-\mathbf{v}_{st}t)$.
13:  end for
14:end for
15:return $Z$.
Algorithm 5 Simulation of a Damped Frozen Field via its Spectrum
1:Spatial grid dimension $n$, number of time steps $T$, advection velocity $\mathbf{v}$.
2:Spatial spectral density $S_{XX}$, persistence parameter $\alpha>0$, damping parameter $\beta\geq 0$.
3:Define padded dimensions $n_{\text{pad}}\leftarrow 2n$ and $T_{\text{pad}}\leftarrow 2T$.
4:Define wavenumber vector $\mathbf{k}_{\text{vec}}$ for a grid of size $n_{\text{pad}}$.
5:Define frequency vector $\omega_{\text{vec}}$ for a grid of size $T_{\text{pad}}$.
6:Initialize complex grid in frequency domain $\widetilde{Z}\in\mathbb{C}^{n_{\text{pad}}\times n_{\text{pad}}\times T_{\text{pad}}}$.
7:for each index $(i,j,t)$ in the padded grid dimensions do
8:  Get corresponding wavevector $\mathbf{k}\leftarrow(\mathbf{k}_{\text{vec}}[i],\mathbf{k}_{\text{vec}}[j])$ and frequency $\omega\leftarrow\omega_{\text{vec}}[t]$.
9:  $S_{XX\_val}\leftarrow S_{XX}(\mathbf{k})$.
10:  $S_{modulator}\leftarrow\left[(\omega+\mathbf{k}^{\top}\mathbf{v})^{2}+(\beta\omega)^{2}\right]^{-\alpha}$.
11:  $S\leftarrow S_{XX\_val}\cdot S_{modulator}$.
12:  Generate complex white noise $\gamma\sim\mathcal{CN}(0,1)$.
13:  $\widetilde{Z}[i,j,t]\leftarrow\sqrt{S}\cdot\gamma$.
14:end for
[…]
16:Take the real part $Z_{\text{pad}}\leftarrow\Re(Z_{\text{pad}})$.
17:Crop to original dimensions $Z\leftarrow Z_{\text{pad}}[0:n-1,0:n-1,0:T-1]$.
18:return $Z$.
Appendix B Derivation of transport-diffusion PDE
In this section we expand on the link between Distributed Frozen Field and the solution of a transport-diffusion PDE.
We recall that a Distributed Frozen Field $Z$ is the convex combination of Frozen Fields $\{Z_{i}\}_{i=1}^{n_{v}}$, i.e.
$Z(\mathbf{s},t)=\sum_{i=1}^{n_{v}}p_{i}Z_{i}(\mathbf{s},t)=\sum_{i=1}^{n_{v}}p_{i}X_{S}\left(\mathbf{s}-\mathbf{v}_{i}t\right),$ (20)
where we called $\mathbf{v}_{i}=\mathbf{M}_{i}\mathbf{v}$, namely the velocity of field $i$, with corresponding weight $p_{i}=p(\mathbf{M}_{i}\mathbf{v})$. By definition, each Frozen Field $Z_{i}$, $i=1,\dots,n_{v}$, satisfies
$\frac{\partial Z_{i}(\mathbf{s},t)}{\partial t}+\mathbf{v}_{i}\cdot\nabla Z_{i}(\mathbf{s},t)=0,$ (21)
with the initial condition $Z_{i}(\mathbf{s},0)=X_{S}(\mathbf{s})$.
Computing the time derivative of $Z(\mathbf{s},t)$ in (20), we get
$\frac{\partial Z(\mathbf{s},t)}{\partial t}=\sum_{i=1}^{n_{v}}p_{i}\frac{\partial Z_{i}(\mathbf{s},t)}{\partial t}=-\sum_{i=1}^{n_{v}}p_{i}\left(\mathbf{v}_{i}\cdot\nabla Z_{i}(\mathbf{s},t)\right),$
where the last step was uptained by plugging in (21). Now, substitute $\mathbf{v}_{i}=\overline{\mathbf{v}}+\Delta\mathbf{v}_{i}$:
$\frac{\partial Z(\mathbf{s},t)}{\partial t}=-\overline{\mathbf{v}}\cdot\left(\sum_{i=1}^{n_{v}}p_{i}\nabla Z_{i}(\mathbf{s},t)\right)-\sum_{i=1}^{n_{v}}p_{i}(\Delta\mathbf{v}_{i}\cdot\nabla Z_{i}(\mathbf{s},t)).$ (22)
Since $\sum_{i=1}^{n_{v}}p_{i}\nabla Z_{i}(\mathbf{s},t)=\nabla Z(\mathbf{s},t)$ thanks to the definition (20), we can rearrange (22) into
$\frac{\partial Z(\mathbf{s},t)}{\partial t}+\overline{\mathbf{v}}\cdot\nabla Z(\mathbf{s},t)=-\sum_{i=1}^{n_{v}}p_{i}(\Delta\mathbf{v}_{i}\cdot\nabla Z_{i}(\mathbf{s},t)).$ (23)
For each component $i=1,\dots,n_{i}$,
$(\nabla Z_{i})\cdot\Delta\mathbf{v}_{i}=\nabla\cdot(Z_{i}\Delta\mathbf{v}_{i})-Z_{i}(\nabla\cdot\Delta\mathbf{v}_{i}),$
and $\nabla\cdot\Delta\mathbf{v}_{i}=0$ since $\Delta\mathbf{v}_{i}$ does not vary with $(\mathbf{s},t)$. Thus, (23) becomes
$\frac{\partial Z(\mathbf{s},t)}{\partial t}+\overline{\mathbf{v}}\cdot\nabla Z(\mathbf{s},t)=-\nabla\cdot\left(\sum_{i=1}^{n_{v}}p_{i}\,Z_{i}(\mathbf{s},t)\Delta\mathbf{v}_{i}\right),$
as shown in the paper.